# arrow size=1pt, every edge=line width=0.5pt

### v1

@@ 5 Summary and outlook
In this paper, we have calculated and analyzed the leading two-loop corrections to the nucleon axial-vector coupling $g_{A}$, that is of fundamental importance in low-energy QCD. We used two covariant versions of baryon chiral perturbation theory, namely we applied the EOMS and the IR renormalization schemes. The pertinent expression for the fourth order corrections $\sim M^{4}$, denoted as $\alpha_{4}$, $\gamma_{4}$ and $\beta_{4}$, as defined in Eq. (1), are explicitly given in Eq. (30), Eq. (4) and Eq. (4), respectively. We have used two sets of the dimension-two and dimension-three LECs from a combined study of the $\pi N\to\pi N$ and $\pi N\to\pi\pi N$ processes in the EOMS scheme and calculated the fourth order corrections as shown in Fig. 3. For set 1 of the LECs, the corrections turned out to be rather small, signaling a good convergence of the chiral expansion of $g_{A}$, whereas for set 2 they are larger but still moderate. However, to finally draw conclusions, the fifth order corrections at order $M_{\pi}^{5}$ need to be worked out and a more detailed uncertainty analysis has to be performed. Such work is in progress.

### v2

@@ 5 Summary and outlook
In this paper, we have calculated and analyzed the leading two-loop corrections to the nucleon axial-vector coupling $g_{A}$, that is of fundamental importance in low-energy QCD. We used two covariant versions of baryon chiral perturbation theory, namely we applied the EOMS and the IR renormalization schemes. The pertinent expression for the fourth order corrections $\sim M^{4}$, denoted as $\alpha_{4}$, $\gamma_{4}$ and $\beta_{4}$, as defined in Eq. (1), are explicitly given in Eq. (30), Eq. (4) and Eq. (4), respectively. We have used two sets of the dimension-two and dimension-three LECs from a combined study of the $\pi N\to\pi N$ and $\pi N\to\pi\pi N$ processes in the EOMS scheme and calculated the fourth order corrections as shown in Fig. 3. For set 1 of the LECs, the corrections turned out to be rather small, signaling a good convergence of the chiral expansion of $g_{A}$, whereas for set 2 they are larger but still moderate. However, to finally draw conclusions, the $M^{5}$ corrections need to be worked out and a more detailed uncertainty analysis has to be performed. Such work is in progress.